OpenStudy (anonymous):

integral of 3x^2 +3x +3/x^2 + 1

OpenStudy (anonymous):

divide first

OpenStudy (anonymous):

I did .. and got 3 + 3x/x^2 + 1

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

the anti derivative of \(3\) is \(3x\)

OpenStudy (anonymous):

\[3\int\frac{xdx}{x^2+1}\] a mental u - sub gives the log

OpenStudy (anonymous):

i.e. \(u=x^2+1, du =2xdx, \frac{du}{2}=xdx\) you get \[\frac{3}{2}\int \frac{du}{u}=\frac{3}{2}\ln(u)=\frac{3}{2}\ln(x^2+1)\]

OpenStudy (anonymous):

im lost .. after I divided I got \[\int\limits 3x + 3x \div x ^{2} + 1\]

OpenStudy (anonymous):

ok maybe there was a mistake somewhere, lets see

OpenStudy (anonymous):

\[\int\limits 3x ( 1 + \frac{ 1 }{ x ^{2}+1}\]

OpenStudy (anonymous):

\[\frac{3x^2+3x+3}{x^2+1}=\frac{3(x^2+x+1)}{x^2+1}\] \[=3\frac{x^2+x+1}{x^2+1}=3\left(\frac{x^2+1}{x^2+1}+\frac{x}{x^2+1}\right)\]

OpenStudy (anonymous):

\[=3+\frac{3x}{x^2+1}\]

OpenStudy (anonymous):

then \[\int(3+\frac{3x}{x^2+1})dx=\int 3dx+\int \frac{3xdx}{x^2+1}\] \[=3x+\frac{3}{2}\ln(x^2+1)\]

OpenStudy (anonymous):

oh .. you pulled out the three in the beginning.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but it makes no difference really the trick is to divide it up in to two parts, one with the numeration of \(x^2+1\) or \(3x^2+3\) to give one term of 3, and the other of \(\frac{3x}{x^2+1}\)

OpenStudy (anonymous):

but how did you get \[\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

nevermind .. i see.